Consider testing Upper H 0 : mu equals 20H0: μ=20 against Upper H Subscript a Baseline : mu less than 20Ha: μ<20 where muμ is the mean number of latex gloves used per week by all hospital​ employees, based on the summary statistics nequals=4444​, x overbarxequals=19.219.2​, and sequals=11.211.2. Compute the​ p-value of the test.
Answer #1
population mean μ= 20
sample mean 'x̄= 19.200
sample size   n= 44.00
sample std deviation s= 11.200
std error 'sx=s/√n= 1.688
test stat t ='(x-μ)*√n/sx= -0.47
p value      = 0.319